Problem 3 please I'm not understanding. Please tell how you got the answer step by step and the equation used ! Thanks
Date 1. Calculate AE for a system undergoing an endothermic process in which 25.3 kJ of heat flows and where 4.6 kJ of work is done on the system. 2. Calculate the work associated with the compression of a gas from 52 L. to 33 L at a constant external pressure of 16 atm. 3. How many grams of ethanol (CH.O; MW 46.068 g/mol) are combusted if 3.00 x 10k energy is released during the combustion? Ethanol is known to release 1.368 x 10 upon combustion. g/mol) are combusted if 3.00 × 10%) of g the combustion? Ethanol is known to release 1.368 × 103 kJ/mol 4. A sample of nickel is heated to 99.8 °C and placed in a calorimeter containing 150.0 g HtO at 23.5°C (sH20 = 4.1 84 J/CCg)). After the metal cools, the final temperature of the nickel and H2O mixture is 25.0 °C. If the mass of the sample of nickel was 23.8 g, what is the specific heat capacity of nickel, swi, assuming no heat was lost to the surroundn 5. Calculate the enthalpy change using Hesss Law for this reaction: 2C)HCH( given the following thermochemical equations: a. C2H2(g) + SO2(g) → 2CO2(g) + H2O() b、 C(s) + O2(g) → CO2(g) c. Fh@ + 2(g) → HaO() AH =-1,299.5 kj AH =-393.5 kj AH =-285.8 kj Tha drcromnositioa of hydrogen peroxide (H:O) was studied and the following data w n of hydrogen peroxide (H O2) was studied and the following data w
Problem 3 please I'm not understanding. Please tel
Answer #1

Combustion of ethanol takes place as:

CH_{3}CH_{2}OH\left ( l \right ) + O_{2}\left ( g \right ) \rightarrow 2 CO_{2}\left ( g \right ) + 3 H_{2}O \left ( l \right ) + Heat\left ( energy \right )

As per the question,

1.368 \times 103 kJ of energy is released by 1 mole of ethanol

\therefore 3.00 \times 103 kJ of energy is released by \frac{1\times 3\times 10^{3}}{1.368\times 10^{3}} = 2.193 moles of ethanol.

Grams of ethanol combusted = moles of ethanol combusted \times molar mass of ethanol = 2.193 \times 46.068 g = 101.027g